OpenStudy (anonymous):

Solve the equation Log(x+7) - logx = 3

OpenStudy (anonymous):

ok in this case you can combine the two logs

OpenStudy (anonymous):

What makes this case different?

OpenStudy (anonymous):

2 logs with same base

OpenStudy (anonymous):

Log x?

OpenStudy (jhannybean):

Remember, \(\log(a) + \log(b) = \log(ab)\) and \(\log(a)-\log(b) = \dfrac{\log(a)}{\log(b)}\)

OpenStudy (anonymous):

What I really hate about algebra is, I get an x and I wonder y.

OpenStudy (jhannybean):

So combine them, @BabyKay

OpenStudy (calculusfunctions):

What @Jhannybean meant was\[\log_{c}a -\log_{c}b =\log_{c}\frac{ a }{ b }\]

OpenStudy (anonymous):

the choices are 3.5, 142.7, 0.0070, and 0.0707.

OpenStudy (calculusfunctions):

I don't care what the choices are, You need to learn how to solve.

OpenStudy (anonymous):

I just wanna catch up on my work b4 the semester ends

OpenStudy (anonymous):

I got the answer anyway....it's 0.0070

OpenStudy (calculusfunctions):

Here's an example:\[\log_{3}(x -2)-\log_{3}(x +1)=2 \]\[\log_{3}\frac{ x -2 }{ x +1 } =2\]Now convert logarithm form to exponential form, to get\[\frac{ x -2 }{ x +1 }=3^{2}\]Now multiply both sides by (x + 1), to get\[x -2=9(x +1)\]\[x -2=9x +9\]Solving for x, we get\[x =-\frac{ 11 }{ 8 }\]Keep in mind the restriction in the problem: x > 2. Hence we have no solution for this example but the answer to your question is correct.